OpenStudy (anonymous):

What are the exact solutions of x2 - 3x - 1 = 0?

OpenStudy (anonymous):

you need to factor this out into (x -a)(x-a)=0 with what i have there can you figure it out?

OpenStudy (anonymous):

I have a real difficult time trying to figure these out.

OpenStudy (anonymous):

ok well i am here to help then.

OpenStudy (anonymous):

Thank you! Please do!

OpenStudy (anonymous):

The LHS does not factor.\[x=\frac{1}{2} \left(3\pm \sqrt{13}\right) \]

OpenStudy (anonymous):

Um, pretty sure you're not going to be able to factor that @PaulK... Quadradic: \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

a = 1 b = -3 c = -1

OpenStudy (anonymous):

ax\(^2\) + bx + c = 0

OpenStudy (anonymous):

well lots of learning done here by giving out of answers so quickly....

OpenStudy (anonymous):

I didn't give out any answer. Only one answer was given out. ;-)

OpenStudy (anonymous):

Thanks guys! I appreciate the help! :D

OpenStudy (anonymous):

The reason you can't factor it is because there is nothing that can be multiplied together to give you -1 and added together to give you -3

OpenStudy (anonymous):

F O I L in other words

OpenStudy (anonymous):

One more question..... the end equations... is it positive 3 or negative 3?

OpenStudy (anonymous):

@thenderson75 Look at the original equation and see if you can identify the coefficients a, b, & c ok?

OpenStudy (anonymous):

I already got the equation and I've figured the whole thing but I wanted to make sure if it was positive or negative before I submit the answer.

OpenStudy (anonymous):

There will be two answers because you have a second-order equation, aka. a quadratic equation. Even if it's the same answer twice, you'll always have two answers.

OpenStudy (anonymous):

Type it in your calculator once with the addition and once with the subtraction and you'll see ;-)

OpenStudy (anonymous):

It becomes some messy decimals

OpenStudy (anonymous):

But if you must know, that's how

OpenStudy (anonymous):

It stops at the equation and one of the choices is negative and one is positive.

OpenStudy (anonymous):

In a lot of real world problems you'll be throwing away the nonsensical negative answer (i.e.: if "x" was "t", and "t" meant "time", you can't have negative time)

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (anonymous):

1/2 (3+\(\sqrt{13}\)) = ? 1/2 (3-\(\sqrt{13})\) = ?

OpenStudy (anonymous):

You can do this even on a basic calculator. If it's for a math class, put down both answer in non-decimal form. If it's for a physics or chemistry class or something, see if there's a root you can get rid of because it's nonsensical and put in decimal form with the correct sig figs and the correct units.

OpenStudy (anonymous):

thx guys @agentx5